# Consistency Models Made Easy

### v1

@@ 6 Limitations
One of the major limitations of ECT is that it requires a dataset to tune DMs to CMs, while recent works developed data-free approaches [14, 46, 74] for diffusion distillation. The distinction between ECT and data-free methods is that ECT learns the consistency condition on a given dataset via the self-teacher, while data-free diffusion distillations acquire knowledge from a frozen diffusion teacher. This feature of ECT can be a potential limitation since the training data of some models are unavailable to the public. However, we hold an optimistic view on tuning CMs using datasets different from pretraining. The data composition and scaling for consistency tuning will be a valuable research direction.
7 Conclusion
We propose a simple yet efficient scheme for training consistency models. First, we derive the differential consistency condition and formulate a loss objective based on this condition. Next, we identify the "curse of consistency" of learning CMs and propose several strategies to mitigate these issues. We find that diffusion models can be considered a special case of consistency models with a loose discretization of the consistency condition. Based on these observations and strategies, we propose a streamlined training recipe called Easy Consistency Tuning (ECT). The resulting models, ECMs, obtained through ECT, unlock state-of-the-art few-step generative capabilities at a minimal tuning cost and are able to benefit from scaling. We hope this work will aid in the wider adoption of consistency models within the research community.
@@ Acknowledgments
Zhengyang Geng and Ashwini Pokle are supported by funding from the Bosch Center for AI. Zico Kolter gratefully acknowledges Bosch funding for the lab.
Appendix A Design Choices for ECT
@@
As training progresses, the mapping function should gradually shrink $\Delta t\to 0$. However, the above parameterization does not achieve this. An alternative parameterization is to exponentially decrease $\Delta t$. We can rewrite the ratio between $r$ and $t$ as:
$\frac{r}{t}=1-\frac{1}{q^{a}}$ (17)
@@
As discussed in Sec. 2, iCT uses pseudo-Huber metric [7] to mitigate the perceptual bias caused by the LPIPS metric [76],
$L({\mathbf{x}},{\mathbf{y}})=\sqrt{\|{\mathbf{x}}-{\mathbf{y}}\|_{2}^{2}+c^{2}}-c,\quad c>0.$ (19)
@@
where we have decomposed the differential of pseudo-Huber loss into an adaptive weighting term and the differential of the squared $L_{2}$ loss. Therefore, we retain the squared $L_{2}$ metric used in DMs, and explore varying adaptive weighting terms.
@@
As noted in Sec. 2, iCT [64] employs a discrete-time curriculum given by  Eq. 8. This curriculum divides the noise horizon $[0,T]$ into $N$ smaller consecutive subintervals to apply the consistency loss, characterized by non-overlapping segments $[t_{i},t_{i+1}]$, and gradually increases the number of intervals $N=10\to 1280$. However, the "boundary" condition of this schedule is to start with the number of intervals to $N=1$, learning a model solely mapping samples at noise levels $T_{\text{max}}$ to the clean data ${\mathbf{x}}_{0}$, largely distinct from the classic diffusion models training. We instead investigate the design principles of a continuous-time schedule whose "boundary" condition yields diffusion pretraining, i.e., constructing training pairs of $r=0$ for all $t$ at the beginning.
Appendix B Exploring the Design Space of Consistency Models
Due to ECT’s efficiency, we can explore the design space of CMs at minimal cost. We specifically examine the weighting function, training schedule, and regularization for CMs.
Our most significant finding is that controlling gradient variances and balancing the gradients across different noise levels are fundamental to CMs’ training dynamics. Leveraging the deep connection between CMs and DMs, we also improve the diffusion pretraining using our insights.
@@ Appendix B Exploring the Design Space of Consistency Models
Forward processes with different noise schedules and model parameterizations can be translated into each other at the cost of varying weighting functions [30]. From our experiments on a wide range of weighting schemes, we learn three key lessons.
@@
$(1)$ There is no free lunch for weighting function, i.e., there is likely no universal regular weighting $\bar{w}(t)$ that can outperform all other candidates for different datasets, models, and target metrics for both 1-step and 2-step generation.
We refer these results to Tab. 2, including $\mathrm{SNR}(t)=\nicefrac{{1}}{{t^{2}}}$, $\mathrm{SNR}(t)+1=\nicefrac{{1}}{{t^{2}}}+1$ [58], EDM weighting $\mathrm{SNR}(t)+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}=\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}$ [26], and Soft-Min-SNR weighting $\nicefrac{{1}}{{(t^{2}+\sigma_{\mathrm{data}}^{2})}}$ [16, 9], where $\mathrm{SNR}(t)=\nicefrac{{1}}{{t^{2}}}$ is the signal-to-noise ratio in our setup.
On CIFAR-10, the weighting $\bar{w}(t)=\nicefrac{{1}}{{(t-r)}}$ from the discretization of consistency condition in Eq. 12 achieves the best 1-step FID, while the square root of $\mathrm{SNR}(t)$, $\bar{w}(t)=\sqrt{\mathrm{SNR}(t)}=\nicefrac{{1}}{{t}}$, produces the best $\text{FD}_{\text{DINOv2}}$. On ImageNet 64$\times$64, considering that we have already had the adaptive weighting $w(\Delta)$, the uniform weighting $\bar{w}(t)\equiv 1$ can demonstrate the best 1-step FID when tuning from EDM2 [28]. In contrast to 1-step FIDs, a wider range of regular weighting $\bar{w}(t)$ produces close 2-step FIDs for ECMs.
When starting on a new dataset with no prior information, $\bar{w}(t)=\mathrm{SNR}(t)+n$ is a generally strong choice as the default regular weighting of data prediction models (x-pred), corresponding to using v-pred [58] or flow matching [38, 41] as model parameterization when $n=1$.
Table 3: Performance of ECMs trained with varying adaptive weighting factor on ImageNet 64$\times$64.
@@
Figure 6: Influence of adaptive weighting $w(\Delta)=\nicefrac{{1}}{{(\|\Delta\|_{2}^{2}+c^{2})^{p}}}$ on pretraining using varying $p$.
Beyond ECT, we further investigate the role of adaptive weighting $w(\Delta)$ in pretraining on a toy SwissRoll dataset using flow matching [38] and a simple MLP network.
@@
As we increase the strength of adaptive weighting, flow models become easier to sample from. Surprisingly, even $p=1$ demonstrates strong few-step sampling results when pretraining the flow model. See Fig. 6 for visualization.
$(3)$ The regular weighting $\bar{w}(t)$ and the adaptive weighting $w(\Delta)$ are compatible. This compatibility is how we build ECMs in this work, using adaptive weighting $w(\Delta)$ to achieve strong 1-step sampling capabilities and regular weighting $\bar{w}(t)$ to improve 2-step sampling.
@@
On CIFAR-10, the constant mapping function with $n(t)\equiv 1$ achieves 1-step FID of 4.06 at $200$k iterations, worse than the 1-step FID of 3.86 by $n(t)=1+\frac{k}{1+e^{bt}}$. Under our forward process (${\mathbf{x}}_{t}={\mathbf{x}}_{0}+t\cdot\epsilon$) and model parameterization (EDM [26]), the constant mapping function also incurs training instability on ImageNet 64$\times$64, likely due to the imbalanced gradient flow.
The role of the mapping function, regarding training, is to balance the difficulty of learning consistency condition across different noise levels, avoiding trivial consistency loss near $t\to 0$. For model parameterizations and forward processes different from ours, for example, flow matching [38, 41, 61], we advise readers to start from the constant mapping function due to its simplicity.
@@
We find that CMs benefit significantly from dropout [20]. On CIFAR-10, we apply a dropout of $0.20$ for models evaluated on FID and a dropout of $0.30$ for models evaluated on $\text{FD}_{\text{DINOv2}}$. On ImageNet 64$\times$64, when increasing the dropout rate from $0.10$ to $0.40$, the 2-step FID decreases from $4.53$ to $3.24$. Increasing the dropout rate further can be helpful for 1-step FID under certain weighting functions, but the 2-step FID starts to deteriorate. In general, we optimize our model configurations for efficient 2-step generation and, therefore, choose the dropout rate of $0.40$ for ECM-S.
@@
Compared with no shrinkage of $\Delta t$, where $\Delta t\approx\mathrm{d}t$ throughout, we find that shrinking $\Delta t\to\mathrm{d}t$ results in improved performance for ECMs. For example, on CIFAR-10, starting ECT directly with $\Delta t\approx\mathrm{d}t$ by setting $q=256$ (corresponding to $\nicefrac{{r}}{{t}}\approx 0.99$) leads to quick improvements in sample quality initially but results in slower convergence later on. The 1-step FID drops from 3.60 to 3.86 using the same 400k training iterations compared to gradually shrinking $\Delta t\to\mathrm{d}t$. On ImageNet 64$\times$64, using $\Delta t\approx\mathrm{d}t$ with $q=256$ from the beginning results in divergence, as the gradient flow is highly imbalanced across noise scales, even when initializing from pretrained diffusion models.
@@
BigGAN [5] 326.66 1
StyleGAN2-ADA [25] 305.92 1
StyleGAN-XL [59] 204.60 1
@@
Training Generative Models in 1 GPU Hour.
To demonstrate ECT’s efficiency, we set up a training budget of one GPU hour. We optimize ECT’s setup to achieve the best 2-step FID within this budget, using $\Delta t\approx\mathrm{d}t$ with $q=256$ and a faster EMA rate of $0.9993$. Through $8000$ gradient descent steps with a batch size of $128$, within 1 hour on a single A100 40G GPU, ECT achieves a 2-step FID of $2.73$, outperforming Consistency Distillation (2-step FID of $2.93$) trained with 800k iterations at batch size 512 and the LPIPS [76] metric.
Pretraining using ECT.
For the largest $\Delta t=t$, ECT falls back to diffusion pretraining with $r=0$ and thus $f_{\theta}({\mathbf{x}}_{r})={\mathbf{x}}_{0}$. We pretrain EDM [26] on the CIFAR-10 dataset using ECT. Instead of using EDM weighting, $\mathrm{SNR}(t)+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}$, we enable the adaptive weighting $w(\Delta)$ with $p=\nicefrac{{1}}{{2}}$ and smoothing factor $c=0$ and a regular weighting $\bar{w}(t)=\nicefrac{{1}}{{t}}$.
Compared with the EDM baseline, ECT brings a convergence acceleration over 2×2$\times$2 × regarding $\text{FD}_{\text{DINOv2}}$, matching EDM’s final performance using less than half of the pretraining budget and largely outperforming it at the full pretraining budget.
EDM pretrained by ECT achieves $\text{FD}_{\text{DINOv2}}$ of $150.39$ for unconditional generation and $121.05$ for class-conditional generation, considerably better than the EDM baseline’s $\text{FD}_{\text{DINOv2}}$ of $168.17$ for unconditional generation and $145.20$ for class-conditional generation, when using the same training budget and inference steps (NFE=35).
@@
Using ECT pretrained models ($\text{FD}_{\text{DINOv2}}$ of $121.05$) and original EDM [26] ($\text{FD}_{\text{DINOv2}}$ of $145.20$), we investigate the influence of pretraining quality on consistency tuning and resulting few-step generative models. Our experiments confirm that better pretraining leads to easier consistency tuning. At the same budget of $204.8\mathrm{M}$ training images, tuning from ECT pretrained models achieves $\text{FD}_{\text{DINOv2}}$ of $128.63$, better than $\text{FD}_{\text{DINOv2}}$ of $152.21$ from EDM.
ECM from the ECT pretraining surpasses SoTA GANs in 1 sampling step and advanced DMs in 2 sampling steps, only slightly falling behind our pretrained models and setting up a new SoTA for modern metric $\text{FD}_{\text{DINOv2}}$. See Tab. 4 for details.
On ImageNet 64$\times$64, ECM-M, initialized from EDM2-M [28], deviates from the power law scaling and achieves better generative performance than the log-linear trend. (See Fig. 4 Right). We speculate that it is due to sufficient pretraining, in which EDM2-M was pretrained by a budget of 2×2$\times$2 × training images of other model sizes (S/L/XL).
Comparison with iCT.
We offer an ablation study to compare ECT and the state-of-the-art iCT [64]. We interpolated back to iCT by removing components from ECT, including dropout rate for 2-step models, continuous-time training schedule through mapping function, and diffusion pretraining. We show the FIDs of 1-step and 2-step sampling and the training iterations in Tab. 5.
Table 5: Comparing ECT and iCT on CIFAR-10.
[…]
ECT 3.86 2.15 200k
w/o Dropout Rate for 2-step Generation 3.56 2.49 200k
w/o Continuous-time Training 3.81 2.87 200k
w/o Diffusion Pretraining ($r=0$) 4.57 3.30 400k
@@
Starting from a proper model size, the benefits from 2-step sampling seem larger than increasing the model size by a factor of two in our setups. In the prior works, iCT [64] employs 2×2$\times$2 × deeper model, but the 1-step generative performance can still be inferior to the 2-step results from ECT. This finding is consistent with recent theoretical analysis [47], which indicates a tighter bound on the sample quality for the 2-step generation compared to the 1-step generation.

### v2

@@ 7 Limitations
One of the major limitations of ECT is that it requires a dataset to tune DMs to CMs. Recent works developed data-free approaches (Gu et al., 2023, Luo et al., 2024, Yin et al., 2023, Zhou et al., 2024a) for diffusion distillation. The distinction between ECT and data-free methods is that ECT learns the consistency condition on a given dataset through the self teacher, while data-free methods transfer knowledge from a frozen diffusion teacher. This feature of ECT can be a potential limitation since the training data of bespoke models are unavailable to the public. However, we hold an optimistic view on tuning CMs using datasets different from pretraining.
In Appendix C, we discuss a data-free variant of Easy Consistency Distillation (ECD), in which we extend the continuous time formulation to Consistency Distillation, by generating self-synthetic data from CM on-the-fly. However, it only applies to the distillation settings. Further data-centric research is needed regarding synthetic data, data composition, and data scaling for consistency models.
Broader Impacts and Ethics Statement
Reproducibility Statement
We provide extensive details of experimental settings and hyperparameters to reproduce our experimental results in Appendix D. We have provided a zip file of our source code in this submission. We plan to release our code to ensure transparency and reproducibility of the results.
@@ Acknowledgments
Zhengyang Geng and Ashwini Pokle are supported by funding from the Bosch Center for AI. Zico Kolter gratefully acknowledges Bosch’s funding for the lab. The FLAME Center at CMU generously supports this research.
We appreciate Yang Song for discussions and feedback. We acknowledge the discussions and comments provided by Tianle Cai, Antony Jia, Diyang Xue, Runtian Zhai, and Yonghao Zhuang during the preparation of the blog post.
Appendix A Motivations behind Design Choices in ECT
@@
As training progresses, the mapping function should gradually shrink $\Delta t\to 0$. However, the above parameterization does not achieve this. An alternative parameterization is to decrease $\Delta t$ exponentially. We assume the ratio between $r$ and $t$ can be written as:
$\frac{r}{t}=1-\frac{1}{q^{a}},$ (17)
@@
As discussed in Sec. 2, iCT uses pseudo-Huber metric (Charbonnier et al., 1997) to mitigate the perceptual bias caused by the LPIPS metric (Zhang et al., 2018),
$L({\mathbf{x}},{\mathbf{y}})=\sqrt{\|{\mathbf{x}}-{\mathbf{y}}\|_{2}^{2}+\epsilon}-\epsilon,\quad\epsilon>0.$ (19)
@@
where we have decomposed the differential of pseudo-Huber loss into an adaptive weighting term and the differential of the squared $L_{2}$ loss. Therefore, we retain the squared $L_{2}$ metric used in DMs, and explore varying adaptive weighting terms which we explore in detail in Appendix B.
@@
As noted in Sec. 2, iCT (Song and Dhariwal, 2023) employs a discrete-time curriculum given by  Eq. 8. This curriculum divides the noise horizon $[0,T]$ into $N$ smaller consecutive subintervals to apply the consistency loss, characterized by non-overlapping segments $[t_{i},t_{i+1}]$, and gradually increases the number of intervals $N=10\to 1280$. However, the "boundary" condition of this schedule is to start with the number of intervals to $N=1$, learning a model solely mapping samples at noise levels $T_{\text{max}}$ to the clean data ${\mathbf{x}}_{0}$, largely distinct from the classic diffusion models training. We instead investigate a continuous-time schedule whose "boundary" condition yields diffusion pretraining, i.e., constructing training pairs of $r=0$ for all $t$ at the beginning.
Appendix B Exploring Design Space & Scaling of Consistency Models
Due to ECT’s efficiency, we can explore the design space of CMs at a minimal cost. We specifically examine the weighting function, training schedule, and regularization for CMs.
Our most significant finding is that controlling gradient variances and balancing the gradients across different noise levels are fundamental to CMs’ training dynamics. Leveraging the deep connection between CMs and DMs, we also improve the diffusion pretraining and the full pretraining+tuning pipeline using our findings.
@@ Appendix B Exploring Design Space & Scaling of Consistency Models
Forward processes with different noise schedules and model parameterizations can be translated into each other at the cost of varying weighting functions (Kingma and Gao, 2024). From our experiments on a wide range of weighting schemes, we learn three key lessons.
@@
$(1)$ There is no free lunch for weighting function, i.e., there is likely no universal timestep weighting $\bar{w}(t)$ that can outperform all other candidates on different datasets, models, and target metrics for both 1-step and 2-step generation.
We refer these results to Tab. 2, including $\mathrm{SNR}(t)=\nicefrac{{1}}{{t^{2}}}$, $\mathrm{SNR}(t)+1=\nicefrac{{1}}{{t^{2}}}+1$ (Salimans and Ho, 2022), EDM weighting $\mathrm{SNR}(t)+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}=\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}$ (Karras et al., 2022), and Soft-Min-SNR weighting $\nicefrac{{1}}{{(t^{2}+\sigma_{\mathrm{data}}^{2})}}$ (Hang et al., 2023, Crowson et al., 2024), where $\mathrm{SNR}(t)=\nicefrac{{1}}{{t^{2}}}$ is the signal-to-noise ratio in our setup.
On CIFAR-10, the weighting $\bar{w}(t)=\nicefrac{{1}}{{(t-r)}}$ from the discretization of consistency condition in Eq. 12 achieves the best 1-step FID, while the square root of $\mathrm{SNR}(t)$, $\bar{w}(t)=\sqrt{\mathrm{SNR}(t)}=\nicefrac{{1}}{{t}}$, produces the best $\text{FD}_{\text{DINOv2}}$. On ImageNet 64$\times$64, considering that we have already had the adaptive weighting $w(\Delta)$, the uniform weighting $\bar{w}(t)\equiv 1$ can demonstrate the best 1-step FID when tuning from EDM2 (Karras et al., 2024). In contrast to 1-step FIDs, a wider range of timestep weighting $\bar{w}(t)$ produces close 2-step FIDs for ECMs.
When starting on a new dataset with no prior information, $\bar{w}(t)=\mathrm{SNR}(t)+n$ is a generally strong choice as the default timestep weighting of data prediction models (x-pred). In this situation, this weighting function corresponds to using v-pred (Salimans and Ho, 2022) or flow matching (Lipman et al., 2022, Liu et al., 2022b) as model parameterization when $n=1$.
Table 3: Performance of ECMs trained with varying adaptive weightings on ImageNet 64$\times$64.
@@
Beyond ECT, we further investigate the role of adaptive weighting $w(\Delta)$ in pretraining on a toy Swiss roll dataset using the parameterization and forward process of flow matching (Lipman et al., 2022) and an MLP network.
@@
As we increase the strength of adaptive weighting, flow models become easier to sample from in a few steps. Surprisingly, even $p=1$ demonstrates strong few-step sampling results when pretraining the flow model. See Fig. 6 for visualization.
@@
Figure 6: Influence of adaptive weighting $w(\Delta)=\nicefrac{{1}}{{(\|\Delta\|_{2}^{2}+\epsilon)^{p}}}$ on pretraining using varying $p$.
On CIFAR-10, the constant mapping function with $n(t)\equiv 1$ achieves 1-step FID of 4.06 at $200$k iterations, worse than the 1-step FID of 3.86 by $n(t)=1+\frac{k}{1+e^{bt}}$. Under our forward process (${\mathbf{x}}_{t}={\mathbf{x}}_{0}+t\cdot\epsilon$) and model parameterization (EDM (Karras et al., 2022)), the constant mapping function incurs training instability on ImageNet 64$\times$64, likely due to the imbalanced gradient flow.
The role of the mapping function, regarding training, is to balance the difficulty of learning consistency condition across different noise levels, avoiding trivial consistency loss near $t\to 0$. For model parameterizations and forward processes different from ours, for example, flow matching (Lipman et al., 2022, Liu et al., 2022b, Sauer et al., 2024), we advise readers to start from the constant mapping function due to its simplicity.
@@
In line with (Song and Dhariwal, 2023), we find that CMs benefit significantly from dropout (Hinton et al., 2012). On CIFAR-10, we apply a dropout of $0.20$ for models evaluated on FID and a dropout of $0.30$ for models evaluated on $\text{FD}_{\text{DINOv2}}$.
On ImageNet 64$\times$64, we note that ECT benefits from a surprisingly high dropout rate. When increasing the dropout rate from $0.10$ to $0.40$, the 2-step FID decreases from $4.53$ to $3.24$. Increasing the dropout rate further can be helpful for 1-step FID under certain timestep weighting $\bar{w}(t)$, but the 2-step FID starts to deteriorate. In general, we optimize our model configurations for 2-step generation and choose the dropout rate of $0.40$ for ECM-S.
@@
Compared with no shrinkage of $\Delta t$, where $\Delta t\approx\mathrm{d}t$ throughout, we find that shrinking $\Delta t\to\mathrm{d}t$ results in improved performance for ECMs. For example, on CIFAR-10, starting ECT directly with $\Delta t\approx\mathrm{d}t$ by setting $q=256$ (corresponding to $\nicefrac{{r}}{{t}}\approx 0.99$) leads to quick improvements in sample quality initially but slower convergence later on. The 1-step FID drops from 3.60 to 3.86 using the same 400k training iterations compared to gradually shrinking $\Delta t\to\mathrm{d}t$. On ImageNet 64$\times$64, $\Delta t\approx\mathrm{d}t$ with $q=256$ from the beginning results in training divergence, as the gradient flow is highly imbalanced across noise levels, even when initializing from pretrained diffusion models.
@@
Training Generative Models in 1 GPU Hour.
To demonstrate the efficiency of ECT and facilitate future studies, we implemented a fast prototyping setting designed to yield results within one hour on a single GPU. This configuration uses a fixed $\Delta t\approx\mathrm{d}t$ by setting $q=256$ in our mapping function (corresponding to $\frac{r}{t}\approx 0.99$), which allows for quick approximation of the differential consistency condition. Through $8000$ gradient descent steps at batch size of $128$, within 1 hour on a single A100 40GB GPU, ECT achieves a 2-step FID of $2.73$, outperforming Consistency Distillation (2-step FID of $2.93$) trained with 800k iters at batch size 512 and LPIPS (Zhang et al., 2018) metric.
Ablation on Pretraining
We conducted a controlled experiment, training CMs from scratch following iCT best practices and tuning EDM using both iCT and ECT schedules. The combined pretraining and fine-tuning cost for ECT is $50\%+6.25\%$ of iCT trained from scratch. Results are presented in Tab. 4.
There is a noticeable performance gap between iCT’s training and ECT’s pretraining+tuning scheme, particularly when evaluating models using modern metrics like $\mathrm{FD}_{\text{DINOv2}}$ (Stein et al., 2024). $\mathrm{FD}_{\text{DINOv2}}$ employs the representation space of the DINOv2-L model (Oquab et al., 2023) to compute distributional distance, which has been shown to better align with human evaluations.
Table 4: Impact of pretraining on model performance
Model $\mathrm{FD}_{\text{DINOv2}}$
iCT 242.30
Pretraining + iCT tuning 200.31
Pretraining + ECT tuning 190.13
EDM 168.16
When scaling up the pretraining+tuning cost to match the overall cost of iCT in class-conditional settings, ECM achieves a $\mathrm{FD}_{\text{DINOv2}}$ of 152.21, significantly outperforming the iCT model trained from scratch (205.11). For context, StyleGAN-XL achieves an $\mathrm{FD}_{\text{DINOv2}}$ of 204.60.
@@
BigGAN (Brock et al., 2018) 326.66 1
StyleGAN2-ADA (Karras et al., 2020) 305.92 1
StyleGAN-XL (Sauer et al., 2022) 204.60 1
@@
Ablation on Tuning Design Choices
Besides quantitatively evaluating the pretraining impact, we conduct further ablation studies regarding fine-tuning design choices. Due to the computational constraints, we tune EDM2-S using iCT’s design choices as the baseline. Results are presented in Tab. 6. These results demonstrate the improvements of the continuous-time training schedule, increased dropout, and weighting functions on ECMs’ generative performance.
Table 6: Ablation study of continuous-time CMs on ImageNet 64$\times$64
Methods 1-step FID 2-step FID
iCT + EDM2 Pretraining 21.09 4.39
+ Continuous-time schedule 14.34 4.33
+ Dropout = 0.40 9.28 3.22
+ $\bar{w}(t)=\nicefrac{{1}}{{t^{2}}}+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}$ 5.51 3.18
Improving Pretraining using Findings in Tuning.
The exploration of the design space through the tuning stage as a proxy led to a question: Can the insights gained during tuning be applied to improve the pretraining stage and, consequently, the entire pretraining+tuning pipeline for CMs? The results of our experiments confirmed this hypothesis.
For the largest $\Delta t=t$, ECT falls back to diffusion pretraining with $r=0$ and thus $f_{\theta}({\mathbf{x}}_{r})={\mathbf{x}}_{0}$. We pretrain EDM (Karras et al., 2022) on the CIFAR-10 dataset using the findings in ECT. Instead of using EDM weighting, $\mathrm{SNR}(t)+\nicefrac{{1}}{{\sigma_{\mathrm{data}}^{2}}}$, we enable the adaptive weighting $w(\Delta)$ with $p=\nicefrac{{1}}{{2}}$ and smoothing factor $c=0$ and a timestep weighting $\bar{w}(t)=\nicefrac{{1}}{{t}}$.
Compared with the EDM baseline, the recipe from ECT brings a convergence acceleration over 2×2$\times$2 × regarding $\text{FD}_{\text{DINOv2}}$, matching EDM’s final performance using less than half of the pretraining budget and largely outperforming it at the full pretraining budget.
EDM pretrained by ECT achieves $\text{FD}_{\text{DINOv2}}$ of $150.39$ for unconditional generation and $121.05$ for class-conditional generation, considerably better than the EDM baseline’s $\text{FD}_{\text{DINOv2}}$ of $168.17$ for unconditional generation and $145.20$ for class-conditional generation, when using the same pretraining budget and inference steps (NFE=35).
@@
Using ECT pretrained models ($\text{FD}_{\text{DINOv2}}$ of $121.05$) and original EDM (Karras et al., 2022) ($\text{FD}_{\text{DINOv2}}$ of $145.20$), we investigate the influence of pretraining quality on consistency tuning and resulting ECMs. Our experiments confirm that better pretraining leads to easier consistency tuning and faster convergence. At the same budget of $204.8\mathrm{M}$ images, tuning from ECT pretrained models achieves $\text{FD}_{\text{DINOv2}}$ of $128.63$, better than $\text{FD}_{\text{DINOv2}}$ of $152.21$ from EDM.
ECM from the ECT pretraining surpasses SoTA GANs in 1 sampling step and advanced DMs in 2 sampling steps, only slightly falling behind our pretrained models and setting up a new SoTA for the modern metric $\text{FD}_{\text{DINOv2}}$. Results can be found in Tab. 5.
On ImageNet 64$\times$64, ECM-M, initialized from EDM2-M (Karras et al., 2024), deviates from the power law scaling and achieves better generative performance than the log-linear trend. (See Fig. 4, Right). We speculate that it is due to a higher pretraining budget, in which EDM2-M was pretrained by 2×2$\times$2 × training images compared with other model sizes (S/L/XL).
@@
In our setups, starting from a proper model size, the improvements from 2-step sampling seem larger than doubling the model size but keeping 1-step sampling. In the prior works, iCT (Song and Dhariwal, 2023) employs 2×2$\times$2 × deeper model, but the 1-step generative performance can be inferior to the 2-step results from ECT. This finding is consistent with recent theoretical analysis (Lyu et al., 2023), which indicates a tighter bound on the sample quality for the 2-step generation compared to the 1-step generation.